OpenStudy (unklerhaukus):

Solve using separation methods:\[\left(y-yx^2\right)\frac{\text dy}{\text dx}=\left(y+1\right)^2\]

OpenStudy (anonymous):

god I hate implicit diffrentiation

OpenStudy (unklerhaukus):

i think it is separable

OpenStudy (unklerhaukus):

\[y\left(1-x^2\right)\frac{\text dy}{\text dx}=\left(y+1\right)^2\]

OpenStudy (unklerhaukus):

\[\frac{y}{\left(y+1\right)^2}{\text dy}=\frac{1}{1-x^2}\text dx\]

OpenStudy (unklerhaukus):

\[\int\frac{y}{\left(y+1\right)^2}{\text dy}=\int\frac{1}{1-x^2}\text dx\]

OpenStudy (unklerhaukus):

now what can i do

OpenStudy (unklerhaukus):

\[\int\frac{y}{\left(y+1\right)\left(y+1\right)}{\text dy}=\int\frac{1}{(1-x)(x+1)}\text dx\]?

OpenStudy (unklerhaukus):

the answer in the back of the book has log's in it

OpenStudy (lgbasallote):

homogeneous!!!! woo!! lol

OpenStudy (anonymous):

\[ \frac{y}{(y+1)^2}=\frac{1}{y+1}-\frac{1}{(y+1)^ 2} \]

OpenStudy (anonymous):

\[ \frac{1}{1-x^2}=\frac{1}{2 (x+1)}-\frac{1}{2 (x-1)} \]

OpenStudy (unklerhaukus):

\[\int\frac{y}{\left(y+1\right)^2}{\text dy}=\int\frac{1}{1-x^2}\text dx\]\[\int\frac{1}{y+1}-\frac{1}{(y+1)^2}=\int\frac{1}{1-x^2}\text dx\]\[\ln|y+1|+\frac 1{y+1}=\int\frac{1}{1-x^2}\text dx\]

OpenStudy (unklerhaukus):

\[\ln|y+1|+\frac 1{y+1}=\int \frac{1}{2 (x+1)}-\frac{1}{2(x-1)} \text dx\] \[\ln|y+1|+\frac 1{y+1}=\frac 12 \left(\ln|x+1|-\ln|x-1|\right)+c\] \[\ln|y+1|+\frac 1{y+1}=\frac 12 \ln\left|\frac{x+1}{x-1}\right|+c\]

OpenStudy (unklerhaukus):

which is the answer in the the back of my book thank you @eliassaab

OpenStudy (unklerhaukus):

is that factorization ment to be obvious

OpenStudy (anonymous):

This is called decomposition in partial fractions.

OpenStudy (unklerhaukus):

ah yes i suppose intermediate steps would have involved A and B which turned out to be 1,1 and 1/2,-1/2

OpenStudy (anonymous):

yes